OpenStudy (anonymous):

\[\int\limits_{5}^{4}x / 4-x ^{2}dx\]

OpenStudy (mimi_x3):

\[\int\limits\frac{x}{4-x^{2}} dx\]?

OpenStudy (mimi_x3):

use \frac{}{} for fractions

OpenStudy (anonymous):

hehe. ok thanks. :)

OpenStudy (anonymous):

can you solve @Mimi_x3 ? :)

OpenStudy (wasiqss):

easiest integral of 2012 ;D

OpenStudy (mimi_x3):

let \(u= 4-x^2\) => \[\frac{du}{-2x} \]

OpenStudy (mimi_x3):

\[\int\limits\frac{x}{4-x^{2}} dx => \int\limits\frac{\cancel{x}}{u} * \frac{du}{-2\cancel{x}} => -\frac{1}{2} \int\limits\frac{1}{u} du\]

OpenStudy (anonymous):

\implies \(\implies\)

OpenStudy (anonymous):

please substitute the value of u @Mimi_x3 . :)

OpenStudy (mimi_x3):

you can do it beatles. just try.

OpenStudy (anonymous):

I can't. I can't do \[\frac{1}{2}\ln \frac{4}{7}\] is the answer. :(

OpenStudy (mimi_x3):

the sub is \(u=4-x^2\) then you just sub it in..

OpenStudy (anonymous):

but the answer on my book is what i type before. :(

OpenStudy (mimi_x3):

wait its a definite integral; just sub in the limits

OpenStudy (wasiqss):

yeh his/her answer is right

OpenStudy (mimi_x3):

\[-\frac{1}{2} \int\limits\frac{1}{u} du => -\frac{1}{2} \ln(u) + C => -\frac{1}{2} \ln(4-x^2) + C\]

OpenStudy (anonymous):

will you do @Mimi_x3 ?

OpenStudy (mimi_x3):

since wasiqs said that its right; then its right lol

OpenStudy (anonymous):

by definite integration. :)

OpenStudy (wasiqss):

i feel honoured lol

OpenStudy (mimi_x3):

wasiqs will do it for you. (:

OpenStudy (anonymous):

@Mimi_x3 your answer is right if it is indefinite integration but in definite will you do it? see me the solution. :)

OpenStudy (anonymous):

change the limits as you do the problem \[u(x)=4-x^2\implies u(5)=4-5^2=-21\] etc

OpenStudy (anonymous):

@Mimi_x3 are you sure that @wasiqss can do that? :)

OpenStudy (anonymous):

then you don't have to convert back, you can leave everything in terms of \(u\)

OpenStudy (anonymous):

ok thank you @satellite73 for your effort but I already know that but my solution can't get the correct answer. :(